OpenStudy (anonymous):

Does the IVP\[(y-7)\frac{dy}{dx}=3x,\]\[y(0)=7\]have a unique solution? I rearranged it to\[\frac{dy}{dx}=\frac{3x}{y-7}\]and differentiated the RHS with respect to 'y' to obtain\[-\frac{3x}{(y-7)^2}.\]Obviously, both functions are not continuous at the point (0,7) [division by zero], and therefore, the IVP has no unique solution. Why did the professor mark me wrong on that?

OpenStudy (anonymous):

i am not sure what an ivp is. but this seems to me like a differentia equation that you would need to separate and integrate to solve.

OpenStudy (anonymous):

why did you not use separation of variables here?

OpenStudy (anonymous):

I don't really have to solve the ODE to determine whether it has a unique solution or not, but it turns out that the aforementioned IVP has two solutions:\[y(x)=7-\sqrt{3}\sqrt{x^2},\]\[y(x)=\sqrt{3}\sqrt{x^2}+7.\]I just don't get why I got it marked wrong.

OpenStudy (anonymous):

I believe it may have been my wording; I made it sound like the IVP has no solution at all. Anyway... thanks for the help!

OpenStudy (anonymous):

I think you have to separate variables first to see if there is any discontinuity in the problem, What do you have to do to view the continuity or domain?

OpenStudy (anonymous):

To check for the existence and uniqueness of a solution to an IVP of a first order ODE, there's a theorem which states that, given\[\frac{dy}{dx}=f(x,y),\]\[y(x_0)=y_0,\]if\[f(x,y)\]and\[\frac{\partial f}{\partial y}\]are both found to be continuous on a region R containing the point\[(x_0,y_0),\]then it's guaranteed to have a unique solution.

OpenStudy (anonymous):

do yes solve for dy/dx and look for breaks in the domain.

OpenStudy (anonymous):

and then take the partial derivative with respect to y of the solution and check for breaks in the domain also...